# Angular control noise in Advanced Virgo and implications for the Einstein Telescope

### 23 Jan 2024 v1

@@ Abstract
With significantly improved sensitivity, the Einstein Telescope (ET), along with other upcoming gravitational wave detectors, will mark the beginning of precision gravitational wave astronomy. However, the pursuit of surpassing current detector capabilities requires careful consideration of technical constraints inherent in existing designs. The significant improvement of ET lies in the low-frequency range, where it anticipates a one million-fold increase in sensitivity compared to current detectors. Angular control noise is a primary limitation for LIGO detectors in this frequency range, originating from the need to maintain optical alignment. Given the expected improvements in ET’s low-frequency range, precise assessment of angular control noise becomes crucial for achieving target sensitivity. To address this, we developed a model of the angular control system of Advanced Virgo, closely matching experimental data and providing a robust foundation for modeling future-generation detectors. Our model, for the first time, enables replication of the measured coupling level between angle and length. Additionally, our findings confirm that Virgo, unlike LIGO, is not constrained by alignment control noise, even under full power operation.
@@ I Introduction
Figure 1: The black line represents the target sensitivity of the Advanced Virgo detector, while the red line indicates the expected level of angular control noise. This curve was calculated based on the circulating power in the cavities of $650$ kW, which is approximately five times higher than that used during O3. On the right, an example of very good agreement is shown between the model, and experimental data is shown. In particular, the plot illustrates the Beam-Spot Motion (BSM) on one of the mirrors. Further clarification is provided later, emphasizing that the BSM is responsible for the angle-to-length coupling.
@@
The Einstein Telescope [1], along with other forthcoming detectors, will advance gravitational wave detection capabilities to an unprecedented level, enabling the observation of previously undetectable events, such as the mergers of intermediate-mass black holes, and conducting precise tests of Einstein’s theory of general relativity under extreme gravitational conditions. Nevertheless, as the objective is to surpass the capabilities of current detectors, it is logical to assume that the technical constraints inherent in existing detectors might hinder the achievement of ET’s goals unless the design is carefully crafted to avoid such limitations. The significant improvement that will be brought about by ET lies primarily in the low-frequency range, where an approximately one million-fold increase in sensitivity is expected compared to current detectors. In this frequency range, the primary limitation of LIGO detectors [2] is due to angular control noise that arises from the need to maintain optical alignment  [3, 4, 5]. Maintaining the operational integrity of the detector requires a high level of alignment stability between the mirrors. Using feedback control ensures that the remaining rotation of the mirrors remains sufficiently low to meet these requirements. However, this comes at the cost of generating control noise that can couple with the detector’s output channel, potentially limiting its performance. Considering the significant expected improvements in this frequency range for ET, it is reasonable to assume that angular control noise may pose a limitation in achieving its target sensitivity. Therefore, it is crucial to be able to evaluate the level of this control noise for ET, and to do this, we needed a precise model. This had not been achieved previously, mainly because there was hope that simplified models would suffice. We took a different approach and developed a more detailed model of the angular control system for the Advanced Virgo detector (AdV) [6]. Our model is closely aligned with experimental data, providing a robust foundation for simulating future-generation detectors. Remarkably, we have successfully replicated, for the first time, the measured coupling level between angle and length; see Fig. 1 and, later in the text, Fig. 7.
@@
Incidentally, our model confirmed that Virgo, unlike LIGO, is not constrained by alignment control noise, as previously shown in [7]. Through the validated model, we show that this limitation would not apply even if the detector was operating at full power; see Fig. 1.
@@
In Section II, we begin by explaining the fundamentals of angular control noise and how mirror rotations can lead to a length signal. We will clarify that assessing the amount of angular noise requires understanding the dynamics of the full control loops. Therefore, we begin the description of angular control loops in Section III, starting with their fundamental component, the plant. Specifically, we present an analytical model to describe the influence of radiation pressure on the suspension system of optics. At the end of the section, we provide a concise overview of how the model was assembled. More details about the plant can be found in Appendix A, while the architecture of the control system can be found in Appendix B. In Section IV, we introduce external disturbances that enter the angular control loops. These disturbances act as a source of angular control noise. For a more detailed description of all external disturbances, see Appendix C. In Section V, we present the level of angular control noise derived from the model and compare it with the experimentally measured noise. A more detailed explanation of the methodology behind generating this noise projection into the detector output can be found in Appendix D. Lastly, we conclude by presenting the same noise projection in the case of the Einstein Telescope in Section VI. In Appendix E, we provide details of the suspension mechanics that were taken into account to produce the estimate of the noise level for ET.
@@ II Angular control noise: the coupling chain
Eq. 2 shows that the cause of the angular noise is the rotation of the optics, which causes the beam to become off-centre. Now, the last thing left to understand is why the optics undergo these movements.
@@ III The opto-mechanical plant and degrees of freedom
Figure 3: Illustration of how a rotation can alter the optical path of a beam if it is offset from the optical axis.
@@
In a gravitational wave detector, a substantial amount of optical power is stored within each arm cavity. This leads to a non-negligible impact of radiation pressure on the mechanics suspending the optics. For each arm cavity, the role of radiation pressure is to couple the motion of the TMs. As explained in Section II, the rotation of one of the TMs offsets the beam on both mirrors. The beam exerts a force on both optics through radiation pressure at a point now offset from the center of rotation, resulting in a torque applied to both TMs. Therefore, the rotation of one TM can induce the motion of the other. On the other hand, in the absence of radiation pressure, the suspensions of each TM operate as mechanically independent systems, with the rotation of one optic not inducing torque on the other optic. When radiation pressure is present, the system cannot be characterized as two independent suspension systems. Instead, it operates as a unified mechanical system, often referred to as an optomechanical system, underscoring the significant role of optics. Fig. 4 illustrates a schematic of the optomechanical system, and its dynamics can be described starting from the equations of a simple torsional pendulum as
@@
Figure 5: Illustration of DoFs under control. $\textnormal{D}_{+}$ stands for Differential Plus, $\textnormal{C}_{+}$ for Common Plus, $\textnormal{N}_{-}$ for North Minus, and $\textnormal{W}_{-}$ for West Minus. Respectively, they include differential tilting of the beam between the arm cavities, common tilting of the beam between the arm cavities, shifting of the beam in the north arm cavity and shifting of the beam in the west arm cavity. We will refer to the pair of DoFs $\textnormal{D}_{+}$ and $\textnormal{C}_{+}$ as “hard modes”, and the pair $\textnormal{N}_{-}$ and $\textnormal{W}_{-}$ as “soft modes”. For more details on the reason of this nomenclature, see Appendix A.
@@ IV External disturbances
From the propagation of noises in the model, we calculated the residual movement of the TMs, which is shown in Appendix C. From the movement of the TMs, it was possible to reconstruct the BSM using Eq. 2. The result is shown in Fig. 7, which shows a measurement of the BMS on one of the TMs and compares it with the model output. It should be noted that this is the first time the measured BSM has been accurately reproduced using a numerical model of the complete optomechanical system.
@@ V Noise projections to the detector output
Figure 8: On the left, the figure displays the level of angular control noise calculated through the model. The plot shows the noise produced by each of the degrees of freedom. On the right are the measured noise levels are shown; this figure has been adapted with permission from [7]. The grey line in both graphs represents the target sensitivity for the O3 run.
@@
Continuing to follow the guidelines outlined starting from Eq. 5, it should be clear that once the movement of the TMs and the BSM is calculated, it is possible to compute the amount of angular control noise coupled to the detector output. However, as previously mentioned, the complexity increases when considering the entire detector, compared to the scenario described by Eq. 5. The angular motion of the TMs coupled to the detector output was calculated using Finesse, and the corresponding result is shown in Fig. 8. More comprehensive details regarding the employed methodology can be found in Appendix D.
The noise projection generated by the model confirms that Advanced Virgo was not limited by angular control noise during O3. This aligns with the experimental measurement of noise levels, as depicted in Fig. 8. To compare the two projections, it should be noted that the two plots reflect different environmental conditions, which define the magnitude of some of the external disturbances considered. For our analysis, external disturbances refer to environmental conditions that led to a relatively significant angle-to-length coupling. Specifically, the Root Mean Square (RMS) of the BSM, which provides an indication of the level of angle-length coupling, was $0.15$ mm. The data in the plot on the right were collected at a different time, which is not specified by the plot’s author. Depending on environmental conditions, the RMS of the BSM was found to reach a minimum value of $0.05$ mm, while on average it settled around $0.10$ mm. This implies that the coupling level can potentially be reduced by up to three times. Consequently, the level of each curve in the model projections can be lower by as much as a factor of three, making the projection levels compatible with those measured experimentally. In addition, what is important to note is that the sequence of contributions per DoF is consistent with that measured experimentally. ${}_{+}$ is the biggest contributor, followed by ${}_{+}$ and finally the soft modes.
@@
Overall, this indicates that the methodology developed for evaluating the residual rotation of the TMs and its subsequent impact on the detector’s output is successful. This establishes a robust foundation for modeling future detector configurations as well as future-generation detectors.
@@
The noise projection was repeated for the case where Virgo was operated at full power. The result is shown in Fig. 1, indicating that Virgo would not be limited by angular noise even in this scenario. To produce this projection, the input noise and the control system architecture were left unchanged. The key change resulting from the increase in power was the impact of the radiation pressure on the mechanical components. The control filters were adapted to the new plant, but they were still designed to ensure a match with the requirements for the residual motion of the TMs.
VI Conclusions for Virgo
@@
VI.1 Implications for the Einstein Telescope
This projection was obtained by directly transferring the AdV ASC system to ET, assuming that almost everything remains unchanged, which is likely not the case. Specifically, we applied the control architecture described in Section V to the ET segment designed for low-frequency measurements, known as ET-Low-Frequency (ET-LF). The optical layout considered is the one implemented in the corresponding Finesse model [18]. As for the mechanics, we considered a suspension inspired by the Virgo superattenuator; more details are provided in Appendix E.
The two elements that require more attention to understand the noise projection are the control filters and the external disturbances. Regarding the control filters, we employed those from AdV and adapted them to the new system. Since the requirements for the residual rotation of the TMs in ET have not yet been defined, the filters have been designed to roughly match the requirements set for Virgo. However, well-defined requirements for ET are essential to determine the loops gain. Without established requirements, the design of the control filters becomes somewhat arbitrary. In simple terms, for the same input noise, a system that requires more extensive control will generate a higher amount of control noise. An example of this effect can be observed in Fig. 6, where we compare the OLTF of ${}_{+}$ with that of ${}_{+}$. Both modes are hard modes and, therefore, share the same plant. As explained in the previous Section III, from a purely mechanical perspective, the arm cavities are identical systems, each with its hard and soft modes. If the TMs in each cavity move in a ${}_{+}$ or ${}_{+}$ manner, there is no mechanical distinction between them. As a result, it might be logical to assume that the same controller could be employed for both DoFs, and thus the OLTF should be identical. However, this is clearly not the case. As explained in Section III, while the common or differential motion of the TMs may be identical from a mechanical point of view, this is not true from an optical point of view. The motion of ${}_{+}$ defines the differential tilt between the arm cavities, affecting the quality of the dark fringe in the output port. Simply put, the less ${}_{+}$ moves, the better the quality of the dark fringe. Consequently, the control loop for this specific DoF requires a higher gain over a wider frequency range, resulting in a different filter design. This filter starts to roll off later and with a gentler slope compared to the filter for ${}_{+}$. As a result, ${}_{+}$ becomes the main source of injected angular control noise at the detector output; you can see this in Fig. 8. This happens despite the fact that ${}_{+}$ has a high-frequency sensing noise almost an order of magnitude lower than that of ${}_{+}$ see Appendix C.
There is one last difference worth highlighting between the noise projection for Virgo and that for ET. For the Einstein Telescope case, the contribution of the soft modes to the projection (N-- and W--) was excluded. In the scheme inherited from Virgo, these degrees of freedom are controlled using optical levers, known to be extremely noisy sensors; see Appendix C. The projection includes only the contributions of the hard modes (D$+$ and C$+$). These degrees of freedom are controlled using QPDs. Notably, these sensors exhibit significantly better noise performance compared to optical levers. The choice to exclude the soft modes from the projection was made to avoid any potential misconception that ET might be limited by angular control noise solely due to the potential use of optical levers. Instead, to accurately assess whether ET will be limited by angular control noise or not, a thorough analysis involving the entire control system is required. This analysis must be based on well-established requirements for the residual motion of various DoFs. These requirements will then inevitably be translated into requirements for the suspensions and external disturbances.
@@ Appendix C External disturbances
Table 3: Measured imbalance levels of actuators and their corresponding values from the model. Actuators are located at a distance of $r=0.3$ m from the rotational axis. The length driving scheme exclusively employs actuators for the end test masses.
@@
The error signals in Fig. 13 capture the motion of the DoFs whenever it exceeds the sensor noise. Conversely, when the actual motion of the DoFs is lower than this noise, the error signals consist mainly of sensing noise shaped by the feedback action of the control loops. This noise is then reintroduced into the system, ultimately causing the mirrors to move. Using the model, the actual motion of the mirrors has been computed for the given noise sources, and the results are shown in Fig. 20. From the motion of the mirrors, we calculated the BSM on the TMs using Eq. 2; the result is shown in Fig. 7. This is the first time the measured BSM has been accurately reproduced using a numerical model of the optomechanical system.

### 04 Mar 2024 v2

@@ Abstract
With significantly improved sensitivity, the Einstein Telescope (ET), along with other upcoming gravitational wave detectors, will mark the beginning of precision gravitational wave astronomy. However, the pursuit of surpassing current detector capabilities requires careful consideration of technical constraints inherent in existing designs. The significant improvement of ET lies in the low-frequency range, where it anticipates a one-million-fold increase in sensitivity compared to current detectors. Angular control noise is a primary limitation for LIGO detectors in this frequency range, originating from the need to maintain optical alignment. Given the expected improvements in ET’s low-frequency range, precise assessment of angular control noise becomes crucial for achieving target sensitivity. To address this, we developed a model of the angular control system of Advanced Virgo, closely matching experimental data and providing a robust foundation for modeling future-generation detectors. Our model, for the first time, enables replication of the measured coupling level between angle and length. Additionally, our findings confirm that Virgo, unlike LIGO, is not constrained by alignment control noise, even if the detector were operating at full power.
@@ I Introduction
Figure 1: Two examples of the excellent agreement between the angular control noise model developed for Advanced Virgo and experimental data. On the left, the level of noise calculated using the model and that measured experimentally are shown. The experimental data are those presented in [1]. The target sensitivity for Advanced Virgo for O3 is shown in black. On the right, the Beam-Spot Motion (BSM) on one of the mirrors is shown. Further clarification is provided later, emphasizing that the BSM is responsible for the angle-to-length coupling.
@@
The Einstein Telescope [2], along with other forthcoming detectors, will advance gravitational wave detection capabilities to an unprecedented level, enabling the observation of previously undetectable events, such as the mergers of intermediate-mass black holes, and conducting precise tests of Einstein’s theory of general relativity under extreme gravitational conditions. Nevertheless, as the objective is to surpass the capabilities of current detectors, it is logical to assume that the technical constraints inherent in existing detectors might hinder the achievement of ET’s goals unless the design is carefully crafted to avoid such limitations. The significant improvement that will be brought about by ET lies primarily in the low-frequency range, where an approximately one million-fold increase in sensitivity is expected compared to current detectors. In this frequency range, the primary limitation of LIGO detectors [3] is due to angular control noise that arises from the need to maintain optical alignment  [4, 5, 6]. Maintaining the operational integrity of the detector requires a high level of alignment stability between the mirrors. Using feedback control ensures that the remaining rotation of the mirrors remains sufficiently low to meet these requirements. However, this comes at the cost of generating control noise that can couple with the detector’s output channel, potentially limiting its performance. Considering the significant expected improvements in this frequency range for ET, it is reasonable to assume that angular control noise may pose a limitation in achieving its target sensitivity. Therefore, it is crucial to be able to evaluate the level of this control noise for ET, and to do this, we needed a precise model. This had not been achieved previously, mainly because there was hope that simplified models would suffice. We took a different approach and developed a more detailed model of the angular control system for the Advanced Virgo detector (AdV) [7]. Our model is closely aligned with experimental data, providing a robust foundation for simulating future-generation detectors. Remarkably, we have successfully replicated, for the first time, the measured coupling level between angle and length; see Fig. 1.
@@
Incidentally, our model confirmed that Virgo, unlike LIGO, is not constrained by alignment control noise, as previously shown in [1]. Through the validated model, we show that this limitation would not apply even if the detector was operating at full power.
@@
In Section II, we begin by explaining the fundamentals of angular control noise and how mirror rotations can lead to a length signal. We will clarify that assessing the amount of angular noise requires understanding the dynamics of the entire control system. Therefore, we begin the description of angular control loops in Section III, starting with their fundamental component, the plant. Specifically, we present an analytical model to describe the influence of radiation pressure on the suspension system of optics. In the same section, we provide a concise overview of how the entire control system model was assembled. More details about the plant can be found in Appendix A, while the architecture of the control system can be found in Appendix B. In Section IV, we introduce the external disturbances that enter the angular control loops. These disturbances act as a source of angular control noise. For a more detailed description of all external disturbances, see Appendix C. In Section V, we briefly explain how we calculated the level of angular control noise coupled to the detector output and show the noise projection in the scenario where the detector was operated at full power. A more detailed explanation of the methodology behind generating noise projections can be found in Appendix D. Lastly, we conclude by presenting the same noise projection in the case of the Einstein Telescope in Section VI. In Appendix E, we provide details of the suspension mechanics that were taken into account to produce the estimate of the noise level for ET.
@@ II Angular control noise: the coupling chain
Figure 3: Illustration in the x-y plane (on the left) and in the z-x plane (on the right) of a beam incident on a mirror, with an offset $d$ with respect to the rotation axis of the optics. The figure aims to show how a mirror rotation can alter the optical path of a beam due to the offset of the beam.
Eq. 2 shows that the cause of angular noise is the rotation of the optics, which causes the beam to become off center. Now, the last thing left to understand is why the optics undergo these movements.
@@ III The opto-mechanical plant and degrees of freedom
In a gravitational wave detector, a substantial amount of optical power is stored within each arm cavity. This leads to a non-negligible impact of radiation pressure on the mechanics suspending the optics. For each arm cavity, the role of radiation pressure is to couple the motion of the TMs. As explained in Section II, the rotation of one of the TMs offsets the beam on both mirrors. The beam exerts a force on both optics through radiation pressure at a point now offset from the center of rotation, resulting in a torque applied to both TMs. Therefore, the rotation of one TM can induce the motion of the other. On the other hand, in the absence of radiation pressure, the suspensions of each TM operate as mechanically independent systems, with the rotation of one optic not inducing torque on the other optic. When there is radiation pressure, the system cannot be characterized as two independent suspension systems. Instead, it operates as a unified mechanical system, often referred to as an optomechanical system, underscoring the significant role of optics. Fig. 4 illustrates a schematic of the optomechanical system, and its dynamics can be described starting from the equations of a simple torsional pendulum as
@@
Figure 5: Illustration of DoFs under control. $\textnormal{D}_{+}$ stands for Differential Plus, $\textnormal{C}_{+}$ for Common Plus, $\textnormal{N}_{-}$ for North Minus, and $\textnormal{W}_{-}$ for West Minus. Respectively, they include differential tilting of the beam between the arm cavities, common tilting of the beam between the arm cavities, shifting of the beam in the north arm cavity, and shifting of the beam in the west arm cavity. We will refer to the pair of DoFs $\textnormal{D}_{+}$ and $\textnormal{C}_{+}$ as “hard modes”, and the pair $\textnormal{N}_{-}$ and $\textnormal{W}_{-}$ as “soft modes”. For more details on the reason for this nomenclature, see Appendix A. The detector requires precise alignment of the optics along both the yaw and pitch axes. Our analysis focuses specifically on the yaw axis.
@@ IV External disturbances
From the propagation of noises in the model, we calculated the residual movement of the TMs, which is shown in Appendix C. From the movement of the TMs, it was possible to reconstruct the BSM using Eq. 2. The result is shown in Fig. 1, which shows a measurement of the BMS on one of the TMs and compares it with the model output. The contribution of each external disturbance to the BSM is shown in the Appendix C. It should be noted that this is the first time the measured BSM has been accurately reproduced using a numerical model of the complete optomechanical system.
@@ V Noise projections to the detector output
Figure 7: The figure shows the level of angular control noise calculated with the model assuming AdV were operated at full power. Even in this case, the detector would not be constrained by this noise. Specifically, this noise projection was calculated based on a circulating power in the cavities of $650$ kW, approximately five times higher than that used during O3. The target sensitivity for AdV during O3 is represented in black.
@@
Continuing to follow the guidelines outlined starting from Eq. 5, it should be clear that once the movement of the TMs and the BSM is calculated, it is possible to compute the amount of angular control noise coupled to the detector output. However, as previously mentioned, the complexity increases when considering the entire detector, compared to the scenario described by Eq. 5. The angular motion of the TMs coupled to the detector output was calculated using Finesse, and the corresponding result is shown in Fig. 1. More comprehensive details on the methodology used can be found in Appendix D. In the appendix, we show the contribution to angular control noise from each controlled DoF. Fig. 1 shows the total level of control noise.
@@
The noise projection generated by the model confirms that Advanced Virgo was not limited by the angular control noise during O3. The model output shows excellent agreement with experimental measurements, indicating the success of the methodology developed for evaluating the residual rotation of the TMs and its subsequent impact on the detector output.This establishes a robust foundation for modeling future detector configurations and future generation detectors.
@@
The noise projection was repeated for the case where Virgo was operated at full power. The result is shown in Fig. 7, indicating that Virgo would not be limited by angular noise even in this scenario. To produce this projection, the input noise and the architecture of the control system were left unchanged. The key change resulting from the increase in power was the impact of the radiation pressure on the mechanical components. The control filters were adapted to the new plant, but they were still designed to ensure a match with the requirements for the residual motion of the TMs.
V.1 Implications for the Einstein Telescope
This projection was obtained by directly transferring the AdV ASC system to ET, assuming that almost everything remains unchanged, which is likely not the case. Specifically, we applied the control architecture described in Section V to the ET segment designed for low-frequency measurements, known as ET-Low-Frequency (ET-LF). The optical layout considered is the one implemented in the corresponding Finesse model [18]. As for the mechanics, we considered a suspension inspired by the Virgo superattenuator; more details are provided in Appendix E. The two elements that require more attention to understand the noise projection are the control filters and external disturbances. Regarding the control filters, we employed those from AdV and adapted them to the new system. Since the requirements for the residual rotation of the TMs in ET have not yet been defined, the filters have been designed to roughly match the requirements set for Virgo. However, well-defined requirements for ET are essential to determine the loops gain. Without established requirements, the design of the control filters becomes somewhat arbitrary. In simple terms, for the same input noise, a system that requires more extensive control will generate a higher amount of control noise. An example of this effect can be observed in Fig. 6, where we compare the OLTF of ${}_{+}$ with that of ${}_{+}$. Both modes are hard modes and, therefore, share the same plant. As explained in the previous Section III, from a purely mechanical perspective, the arm cavities are identical systems, each with its hard and soft modes. If the TMs in each cavity move in a ${}_{+}$ or ${}_{+}$ manner, there is no mechanical distinction between them. As a result, it might be logical to assume that the same controller could be employed for both DoFs, and thus the OLTF should be identical. However, this is clearly not the case. As explained in Section III, while the common or differential motion of the TMs may be identical from a mechanical point of view, this is not true from an optical point of view. The motion of ${}_{+}$ defines the differential tilt between the arm cavities, affecting the quality of the dark fringe in the output port. Simply put, the less ${}_{+}$ moves, the better the quality of the dark fringe. Consequently, the control loop for this specific DoF requires a higher gain over a wider frequency range, resulting in a different filter design. This filter starts to roll off later and with a gentler slope compared to the filter for ${}_{+}$. As a result, ${}_{+}$ becomes the main source of injected angular control noise at the detector output - see Appendix D. This happens despite the fact that ${}_{+}$ has a high-frequency sensing noise almost an order of magnitude lower than that of ${}_{+}$ see Appendix C.
There is one last difference worth highlighting between the noise projection for Virgo and that for ET. For the Einstein Telescope case, the contribution of the soft modes to the projection (N-- and W--) was excluded. In the scheme inherited from Virgo, these degrees of freedom are controlled using optical levers, known to be extremely noisy sensors; see Appendix C. The projection includes only the contributions of the hard modes (D$+$ and C$+$). These degrees of freedom are controlled using QPDs. Notably, these sensors exhibit significantly better noise performance compared to optical levers. The choice to exclude the soft modes from the projection was made to avoid any potential misconception that ET might be limited by angular control noise solely due to the potential use of optical levers. Instead, to accurately assess whether ET will be limited by angular control noise or not, a thorough analysis involving the entire control system is required. This analysis must be based on well-established requirements for the residual motion of various DoFs. These requirements will then inevitably be translated into requirements for the mechanics and external disturbances.
VI Conclusions
@@
Finally, we have shown that angular control noise has the potential to obstruct ET goals. We have explained how formulating control requirements is crucial for accurately assessing whether the ET will be limited by angular control noise or not.
@@ Appendix C External disturbances
Table 3: Measured imbalance levels of actuators and their corresponding values from the model. The actuators are located at a distance of $r=0.3$ m from the rotational axis. The length driving scheme exclusively employs actuators for the end test masses.
@@
The error signals in Fig. 12 capture the motion of the DoFs whenever it exceeds the sensor noise. Conversely, when the actual motion of the DoFs is lower than this noise, the error signals consist mainly of sensing noise shaped by the feedback action of the control loops. This noise is then reintroduced into the system, ultimately causing the mirrors to move. Using the model, the actual motion of the mirrors was computed for the given noise sources, and the results are shown in Fig. 19. From the motion of the mirrors, we calculated the BSM on the TMs using Eq. 2; the result is shown in Fig. 20. This is the first time the measured BSM has been accurately reproduced using a numerical model of the optomechanical system.